# Mutatis Mutandis: Revisiting the Comparator in Discrimination Testing

### v2

@@ 1.2. Relevant Related Work
The literature on discrimination testing is vast, multidisciplinary, and regional. We refer to (Chen et al., 2024) for a recent survey on testing fairness of ML models, and to (Hutchinson and Mitchell, 2019; Romei and Ruggieri, 2014) for surveys with a multidisciplinary perspective. For a general overview of bias and fairness in AI, see (Álvarez et al., 2024; Ntoutsi et al., 2020; Mehrabi et al., 2022). Throughout the paper, we consider individual discrimination. Namely, we are interested at testing whether an individual instance of a protected group has been discriminated. The notion of group discrimination tackles instead discrimination against a subset of individuals from a protected group. However, Binns (2020) argues that the apparent conflict between individual and group fairness is an artifact of fairness metrics rather than reflecting different normative principles.
@@
In this subsection, we position our work relative to three recent papers that we view as directly applicable to the modeling problem of revisiting the comparator. To the best of our knowledge, we are the first to explicitly tackle this problem. Later in Section 5, we provide a representative survey of discrimination testing tools, with a focus on the comparator chosen by those tools. We do so as we first want to introduce the CP and MM comparators (Sections 2 and 3), and illustrate their impact on a discrimination testing problem (Section 4).
@@
First, Kohler-Hausmann (2019) argues that discrimination testing focuses on answering counterfactual questions about the complainant. She refers to the current approach as the causal counterfactual model of discrimination, and states that it is flawed due to its treatment of protected attributes. In particular, using a constructivist approach to race, she argues that comparing similar individuals that only differ on skin color fails to capture how race affects all other aspects of these individuals’ lives. We draw considerably from Kohler-Hausmann (2019) as she writes the first critique to the CP comparator. However, different from Kohler-Hausmann (2019), we propose an alternative to the CP comparator in the form of the MM comparator and embrace the role of counterfactual reasoning within discrimination testing. This last point is important as Kohler-Hausmann’s view (see, e.g., Hu and Kohler-Hausmann (2020)) is largely against causal modeling for discrimination testing and is suspicious toward attempts to measure the so-called causal effect of race. She fails, though, to provide an alternative to the current model. We disagree with her position toward causal modeling. What we agree with her is that we can (and, potentially, should) do better than the CP comparator. The MM comparator is a response to Kohler-Hausmann (2019).
@@
Second, Wachter et al. (2020) state that non-discrimination law aims for substantive equality, meaning it should aim at changing the status quo of the societal context of interest. [^3] Their argument is specific to the European Union. Although it seems not to be the case for US non-discrimination law (Barocas and Selbst, 2016), the goal of substantive equality is implicit in many of the fair ML definitions used by US researchers as argued by Wachter et al. (2020). Under this premise, they classify fair ML methods into bias preserving and bias transforming based on whether they assume (the former) or not (the latter) a neutral status quo. Although Wachter et al. (2020) make a strong case for ML methods as societal shapers, they lack clarity on how these methods are to be implemented, specially, with respect to non-discrimination law. We tackle this gap by formalizing the MM comparator. Our reading of Wachter et al. (2020) is that non-discrimination law must not aim at reaching equality of outcomes only, but also at imposing equality of opportunities in a systematic way. Such a step requires envisioning an ideal target, which needs to be generated, for the current context to aim at. The MM comparator represents the goal of substantive equality in the context of discrimination testing.
@@
Third, Álvarez and Ruggieri (2023) introduce a similar classification to the CP and MM comparators distinction. Drawing from Kohler-Hausmann (2019)’s discussion on the causal counterfactual model of discrimination, they state that discrimination testing tools can be broadly classified into comparators that do and do not consider the downstream effects of the protected attribute on all other attributes. They offer a valuable first step to the goal of revisiting the comparator, but fall short in formalizing their claims on the comparator. They neither define these kinds of comparators nor they survey the quantitative tools to illustrate them. On top of addressing these shortcoming from Álvarez and Ruggieri (2023), we discuss the role of counterfactual reasoning in discrimination testing, which they take for granted.
@@
We note that the causal treatment of discrimination, specifically, using SCM is a common practice. See (Makhlouf et al., 2024) for a survey. In particular, Kilbertus et al. (2017) are the first to formulate problems like direct, indirect, and proxy discrimination using SCM, while Plecko and Bareinboim (2024) try to summarize discrimination within a general causal fairness model using SCM. Works like these are relevant, but lack interdisciplinary depth. For instance, both papers fail to discuss non-discrimination law and how exactly the modeling tools based on their formulations can be (or need to be) used as evidence. Their treatment of the subject is detached from the real-world application. These papers, e.g., do not discuss the comparator. To be clear, such papers have a different scope—mainly to formulate discrimination using SCM—but, because of such approach, fail to address the role of causality in discrimination testing. [^4] Such a discussion is more explicit in the social sciences, in which causal models are also used for testing discrimination. See, e.g., Heckman (1998). In that sense, our work complements these and similar works by revisiting the comparator.
@@ 2. The Comparator as a Causal Modeling Problem
Testing for discrimination amounts to a two-step modeling problem centered on deriving the comparator $i^{\prime}$, and comparing it to the discrimination claim’s complainant $i$. In this section, we focus on the derivation of the comparator, motivating the need for $i^{\prime}$ and its inherent challenges. Importantly, we argue that the derivation of the comparator is a causal modeling problem not only because we use the comparator to demonstrate whether the decision is or not caused by the protected attribute, but also because $i^{\prime}$ is a counterfactual representation of the factual $i$.
@@ 2.1. The Setup
Non-discrimination law mandates that a decision-maker—be it algorithmic or human—treats similar individuals similarly regardless of the protected attribute (Kleinberg et al., 2019). Discrimination testing tools uncover discrimination by finding and comparing similar individuals to isolate the effect of the protected attribute on the decision outcome (Romei and Ruggieri, 2014). Formally, letting $A$ represent the protected attribute, $\mathbf{X}$ the set of non-protected attributes, and $Y$ the decision outcome, we write this two-step modeling problem into Definitions 2.1 and 2.2, respectively.
@@
(The CP Comparator) Let the individual profile $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ represent the complainant $i$ on which the discrimination claim is based. We define a ceteris paribus (CP) comparator $i^{\prime}$ as an individual profile $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$ such that:
@@
where $d$ is a distance function [^5] The issue of engineering fairness is, without doubts, challenging (Scantamburlo, 2021). In this regard, we believe that the definition of the distance function $d$ is not only a technical choice (see the survey (Blanco-Mallo et al., 2023)), but is necessarily specific of the application domain at hand. For instance, in job screening, $d$ can result from a weighting of the skills of job candidates: how to set the weights is a non-trivial issue, which requires an agreement among multiple stakeholders. For this reason, in this paper, we make no assumption/preference about the definition of $d$. (the closer to 00 the more similar are $\mathbf{x}_{i}$ and $\mathbf{x}_{i^{\prime}}$) and ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ a threshold.
Importantly, it is possible to provide a hypothetical comparator (Weerts et al., 2023), meaning we may find or generate an $i^{\prime}$ similar under $d$ to $i$. The comparator derived from (2) is the standard comparator in discrimination testing (Heckman, 1998; Kohler-Hausmann, 2019). [^6] There is a conception of discrimination testing that is non-comparative, though it is not widely used. See Binns (2020). We refer to it as the ceteris paribus, or “with all else being equal,” (CP) comparator as, ideally, we expect to find an $i^{\prime}$ profile all equal but for the protected attribute to the $i$ profile when $\epsilon=0$. Intuitively, it operationalizes the basic axiom underpinning non-discrimination law—that of treating similar individuals similarly—especially when the decision-maker $f$ is, realistically, only allowed to use the set of non-protected attributes when making a decision, meaning $Y=f(\mathbf{X})$. In Section 3 we come back to the CP comparator and revisit it explicitly.
@@
(Difference in Outcomes) Given the individual profiles $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ and $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$, we test the discrimination claim by focusing on the difference in outcomes:
@@
where, $\delta$ may, e.g., be greater or equal to 0 or need to follow a known distribution depending on the legal context.
The difference in outcomes (3) is not based on a literal comparison, meaning $y_{i^{\prime}}-y_{i}$, but on a difference in probabilities. It implies the need to construct some measure of uncertainty (hence, why we write $\delta$ in terms of probabilities) around the decision outcome for each individual profile. Formally, $\delta$ is the difference between two point estimates that requires the use of a statistical estimator. To obtain it, therefore, we need either multiple observations centered on these two profiles or strong assumptions about the behavior of these two profiles. In fact, it seems that legal experts require or, at least, expect this kind of comparison (FRA, 2018). [^7] We believe that this practice should not come as a surprise to the reader: if it takes more than one throw to check whether a coin is fair, why should it not be the same for a serious accusation like discrimination? Intuitively, we wish to control for chance from (or rule out out uncertainty in) the decision process in question to have (some) certainty on the pattern we are testing for. This is not possible under a literal comparison.
Together, these two definitions describe the general setup implicit to all (individual) discrimination tools. We provide an overview of these tools in Section 5. Given our focus on revisiting the standard comparator, in this work we focus on Definition 2.1. Clearly, whatever we derive as the comparator in (2) conditions what outcomes are compared in (3) and, in turn, what we test for and uncover as discrimination when using any discrimination testing tool.
@@ 2.2. On (Establishing) Discrimination
The comparator is intrinsic to how we conceive discrimination. This conception is comparative (Westen, 1982; Weerts et al., 2023; Kohler-Hausmann, 2019). We draw a distinction between describing versus proving something as discriminatory, which helps understand discrimination as a social phenomenon (the former) and as a modeling problem (the latter). Let us start with describing discrimination. Quoting from Weerts et al. (2023), who base their definition on Lippert-Rasmussen (2006): “discrimination can generally be characterized by the morally objectionable practice of subjecting a person (or group of persons) to a treatment in some social dimension that, for no good reason, is disadvantageous compared to the treatment awarded to other persons who are in a similar situation, but who belong to another socially salient group.” A group is considered to be socially salient “if perceived membership of it is important to the structure of social interactions across a wide range of social contexts” (Weerts et al., 2023). Defining the socially salient group, e.g., can only be done by agreeing on a specific, shared social context. Arriving to this point, e.g., by recognizing certain groups under non-discrimination law, captures the social phenomenon.
@@
At the core of this view on discrimination there is a clear comparative element. We are not interested in the difference in treatment between any two individuals from different socially salient groups, but in the difference in treatment between two similar (or similarly situated) individuals from different socially salient groups. Hence, discrimination occurs when similar (or similarly situated) individuals that differ on membership to a socially salient group are treated differently. Under this setting, discrimination becomes the opposite of equality (Weerts et al., 2023). Also under this setting, the comparator becomes a manifestation on what similar (or similarly situated) individuals look like.
@@
Let us now address proving discrimination. We consider the European Union (EU) context because, one, most fair ML works consider only the US context and, second, the EU—with the AI Act and GDPR—currently leads on the regulation front for algorithms (Álvarez et al., 2024). In doing so, we want to show the complex pipeline behind establishing discrimination, and emphasize that it does not begin nor end with Section 2.1. Beyond that, although the discussion below concerns EU non-discrimination law, the modeling problem formalized in Section 2.1 is shared among most (Western) conceptions of justice (Westen, 1982) and it is, thus, a robust problem to the legal context. There are two forms of discrimination conceived by EU non-discrimination law: direct and indirect discrimination. Under direct discrimination, the decision-maker uses the protected attribute for the decision, while under indirect discrimination the decision-maker uses a supposedly neutral attribute that is informative of the protected attribute. [^8] In US non-discrimination law, direct discrimination is refer to as disparate treatment and indirect discrimination as disparate impact (Barocas and Selbst, 2016). The EU context almost fully translates onto the US context, especially when it comes to direct discrimination and disparate treatment. For our purposes, the goal of substantive equality over formal equality is specific to the EU (Wachter et al., 2020). We come back to this point in Section 3, but, regarding the overall problem of the comparator, we can generalize across the two legal contexts. The same seems to hold for UK non-discrimination law (Sargeant and Magnusson, 2024). There are four elements in a discrimination claim, which we list below in order (Weerts et al., 2023). We add in italics the implications to modern decision flows.
@@
  • …“a protected characteristic.” The claim must be based on an attribute protected by non-discrimination law. This element, on top of the current discussions on intersectional versus multiple discrimination (Xenidis, 2020), is ongoing as there is a fear that models are creating new protected attributes.
@@
The third element is the one that concerns the comparator as it requires the use of the discrimination testing tools for obtaining the prima facie evidence. Proving discrimination, similar to describing discrimination, has a clear comparative element; here, though, it has a modeling angle because the goal of obtaining evidence is explicit. The idea is still to compare similar (or similarly situated) individuals, though explicitly by defining $d$ in Definition 2.1, materializing what similarity between individuals is within our context.
Defining similarity is precisely the main challenge inherent to modeling the comparator $i^{\prime}$ with respect to the complainant $i$. The social context that describes discrimination can be and often is taken for granted when modeling the comparator as it is well understood. In Example 1.1, e.g., we did not need to motivate females and males as socially salient groups. The same occurs with most fair ML papers that simply list protected attributes like race and gender to which the proposed methods apply (Ruggieri et al., 2023). However, a well-understood context still does not release us from making normative statements that are socially charge when we define similarity between individuals through $d$. We cannot escape defining $d$ either as only through $d$ we derive the comparator $i^{\prime}$ for the complainant $i$. In principle, as long as some suitable $d$ exists, the modeling problem of proving discrimination appears simple and even intuitive. The simplicity, however, breaks down precisely because we must first define similarity between individuals to then obtain similar individuals and, in turn, prove discrimination. In The Empty Idea of Equality, Westen (1982) points out at this limitation of non-discrimination law by arguing that equality is a circular concept: to argue for equality we first must define what it means to be equal. It follows, thus, that similarity is a circular concept. The circularity of similarity matters because, although a discrimination tool’s main goal is to find evidence, which on paper might seem an objective goal, any discrimination testing tool—be it implicitly or explicitly—needs to implement a normative statement on similarity in the form of some $d$ in order to be operational. It is a difficulty inherent to deriving the comparator, be it the standard one or not. Hence, none of these tools can be sold as unbiased in practice.

### v3

@@ 1.2. Related Work
The literature on discrimination testing is vast, multidisciplinary, and regional. We refer to Chen et al. (2024) for a recent survey on testing fairness of ML models, and to Hutchinson and Mitchell (2019) and Romei and Ruggieri (2014) for surveys with a multidisciplinary perspective. For a general overview of bias and fairness in AI, see Álvarez et al. (2024); Ntoutsi et al. (2020) and Mehrabi et al. (2022).
@@
Here, we wish to position our work with respect to three papers that are directly applicable to the modeling problem of revisiting the comparator in discrimination testing. To the best of our knowledge, we are the first to explicitly tackle this problem. Later in Section 5, we provide a selective survey of discrimination testing tools, with a focus on the comparator chosen by those tools. We do so as we first want to introduce the CP and MM comparators in Sections 2 and 3, and illustrate their impact on a discrimination testing problem in Section 4.
@@
First, Kohler-Hausmann (2019) argues that discrimination testing focuses on answering counterfactual questions about the complainant. The work refers to the current approach as the causal counterfactual model of discrimination, and states that it is flawed due to its treatment of protected attributes. In particular, using a constructivist approach to race, the work argues that comparing similar individuals that only differ on skin color fails to capture how race affects all other aspects of these individuals’ lives. We draw considerably from Kohler-Hausmann (2019) as it represents the first critique to the CP comparator. However, different from Kohler-Hausmann (2019), we propose an alternative to the CP comparator in the form of the MM comparator and embrace the role of counterfactual reasoning within discrimination testing. This last point is important as Kohler-Hausmann’s view (see, for instance, Hu and Kohler-Hausmann (2020)) is largely against causal modeling for discrimination testing and is suspicious toward attempts to measure the so-called causal effect of race. Kohler-Hausmann fails, though, to provide an alternative to the current model. We disagree with Kohler-Hausmann (2019)’s position toward causal modeling, and agree that we can and, potentially, should do better than the CP comparator. The MM comparator is our response to Kohler-Hausmann (2019).
@@
Second, Wachter et al. (2020) states that non-discrimination law aims for substantive equality, meaning it should aim at changing the status quo of the societal context of interest. [^5] The argument is specific to the European Union. Although it seems not to be the case for US non-discrimination law (Barocas and Selbst, 2016), the goal of substantive equality is implicit in many of the fair ML definitions used by US researchers as argued by Wachter et al. (2020). Under this premise, the work classifies fair ML methods into bias preserving and bias transforming based on whether they assume (the former) or not (the latter) a neutral status quo. Although Wachter et al. (2020) makes a strong case for ML methods as societal shapers, the work lacks clarity on how these methods are to be implemented, specially, with respect to non-discrimination law. We tackle this gap by formalizing the MM comparator. Our reading of Wachter et al. (2020) is that non-discrimination law must not aim at reaching equality of outcomes only, but also at imposing equality of opportunities in a systematic way. Such a step requires envisioning an ideal target, which needs to be generated for the current context to aim at. The MM comparator represents the goal of substantive equality in the context of discrimination testing.
@@
Third, Álvarez and Ruggieri (2023) introduces the CP and MM classification but briefly and without the necessary discussion.  [^6] The same criticism applies to the journal version: Álvarez and Ruggieri (2025). We do this on purpose. The CP versus MM discussion merits its own work as it applies to all discrimination testing tools. This is because the focus of the paper is on formulating the discrimination testing method of counterfactual situation testing, which is a possible implementation of the MM comparator. Drawing from Kohler-Hausmann (2019)’s discussion on the causal counterfactual model of discrimination, the work states that discrimination testing tools can be broadly classified into comparators that do and do not consider the downstream effects of the protected attribute on all other attributes. The work offers a valuable first step to the goal of revisiting the comparator, but falls short in formalizing the claims on the comparator. The work neither defines these kinds of comparators nor surveys the quantitative tools to illustrate them. On top of addressing these shortcoming from Álvarez and Ruggieri (2023), we discuss the role of counterfactual reasoning in discrimination testing, which the works takes for granted.
@@
We note that the causal treatment of discrimination, specifically, using SCM is a common practice. See Makhlouf et al. (2024) for a survey. Kilbertus et al. (2017) are the first to formulate problems like direct, indirect, and proxy discrimination using SCM, while Plecko and Bareinboim (2024) try to summarize discrimination within a general causal fairness model using SCM. Works like these are relevant, but lack interdisciplinary depth. For instance, both papers fail to discuss non-discrimination law and how exactly the modeling tools based on their formulations can be or need to be used as evidence. Their treatment of the subject is detached from the real-world application. These papers do not discuss the comparator. To be clear, such papers have a different scope—mainly to formulate discrimination using SCM—but, because of such approach, fail to address the role of causality in discrimination testing. Such a discussion is more explicit in the social sciences, in which causal models are also used for testing discrimination. See, for instance, Heckman (1998). Our work complements these and similar works by revisiting the comparator.
@@ 2. The Comparator as a Causal Modeling Problem
In this section, we argue that testing for discrimination amounts to a two-step modeling problem centered on deriving the comparator $i^{\prime}$ and comparing it to the discrimination claim’s complainant $i$. We argue for the derivation of the comparator, motivating the need for $i^{\prime}$. We also argue that the derivation of the comparator is a causal modeling problem not only because we use $i^{\prime}$ to demonstrate whether the decision is or not caused by the protected attribute, but also because $i^{\prime}$ is a counterfactual representation of the factual $i$.
@@ 2.1. The Setup
Non-discrimination law mandates that a decision-maker, be it algorithmic or human, treats similar individuals similarly regardless of the protected attribute (Kleinberg et al., 2019). Discrimination testing tools uncover discrimination by finding and comparing similar individuals to isolate the effect of the protected attribute on the decision outcome (Romei and Ruggieri, 2014). Formally, letting $A$ represent the protected attribute, $\mathbf{X}$ the set of non-protected attributes, and $Y$ the decision outcome, we write this two-step modeling problem into Definitions 2.1 and 2.2, respectively.
@@
(The CP Comparator) Let the real individual profile $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ represent the complainant $i$ on which the discrimination claim is based on. We define the ceteris paribus (CP) comparator $i^{\prime}$ as a real or fictitious individual profile $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$ such that:
@@
where $d(\cdot,\cdot)$ is a distance function [^7] The issue of engineering fairness is, without doubts, challenging (Scantamburlo, 2021). We believe that the definition of the distance function $d(\cdot,\cdot)$ is not only a technical choice (see the survey by Blanco-Mallo et al. (2023)), but is necessarily specific of the application domain at hand. For instance, in job screening, $d(\cdot,\cdot)$ can result from a weighting of the skills of job candidates: how to set the weights is a non-trivial issue, which requires an agreement among multiple stakeholders. For this reason, in this paper, we make no assumption/preference about the definition of $d(\cdot,\cdot)$. and $\epsilon\in\mathbb{R}^{+}$ a threshold. The closer $d(\cdot,\cdot)$ is to 0 the more similar are $\mathbf{x}_{i}$ and $\mathbf{x}_{i^{\prime}}$.
We emphasize that it is possible to provide a hypothetical comparator (Weerts et al., 2023), meaning we may find or generate an $i^{\prime}$ similar under $d$ to $i$. The comparator derived from Equation (2) is the standard comparator in discrimination testing (Heckman, 1998; Kohler-Hausmann, 2019). [^8] There is a conception of discrimination testing that is non-comparative, though it is not widely used (Binns, 2020). We refer to it as the ceteris paribus, or “with all else being equal,” (CP) comparator as, ideally, we expect to find an $i^{\prime}$ profile all equal but for the protected attribute to the $i$ profile when $\epsilon=0$. Intuitively, it operationalizes the basic axiom underpinning non-discrimination law—that of treating similar individuals similarly—especially when the decision-maker $f$ is, realistically, only allowed to use the set of non-protected attributes when making a decision, meaning $Y=f(\mathbf{X})$. In Section 3 we come back to the CP comparator and revisit it explicitly.
@@
(Difference in Outcomes) Given the individual profiles $\langle\mathbf{x}_{i},a_{i},y_{i}\rangle$ and $\langle\mathbf{x}_{i^{\prime}},a_{i^{\prime}},y_{i^{\prime}}\rangle$, we test the discrimination claim by focusing on the difference in a same decision outcome:
@@
where, $\delta$ may, for example, be greater or equal to 0 or need to follow a known distribution depending on the legal context. Notably, we focus on the same decision between the complainant and the comparator, meaning $y_{i^{\prime}}=y_{i}$.
The difference in outcomes (3) is not based on a literal comparison, meaning $y_{i^{\prime}}-y_{i}$, but on a difference in probabilities. It implies the need to construct some measure of uncertainty (hence, why we write $\delta$ in terms of probabilities) around the decision outcome for each individual profile. Formally, $\delta$ is the difference between two point estimates that requires the use of a statistical estimator. To obtain it, therefore, we need either multiple observations centered on these two profiles or strong assumptions about the behavior of these two profiles. Legal experts require or, at least, expect this kind of comparison (FRA, 2018).We believe that this practice should not come as a surprise to the reader. For instance, if it takes more than one throw to check whether a coin is fair, why should it not be the same for a serious accusation like discrimination? Intuitively, we wish to control for chance from (or rule out out uncertainty in) the decision process in question to have some certainty on the pattern we are testing for. This is not possible under a literal comparison.
Together, these two definitions describe the general setup implicit to all discrimination testing tools. Given our focus on revisiting the standard CP comparator, in this work we focus on Definition 2.1. Clearly, whatever we derive as the comparator following Equation (2) conditions what outcomes are compared in Equation (3) and, in turn, what we test for and uncover as discrimination when using a discrimination testing tool.
@@ 2.2. On (Establishing) Discrimination
The comparator is intrinsic to how we conceive discrimination, and this conception is comparative (Westen, 1982; Weerts et al., 2023; Kohler-Hausmann, 2019). Now, overall, discrimination in practice can be both a familiar and non-trivial concept. We draw a distinction between describing versus proving something as discriminatory, which helps understand discrimination as a social phenomenon (the former) and as a modeling problem (the latter). Let us start with describing discrimination. Quoting from Weerts et al. (2023), who base their own definition on Lippert-Rasmussen (2006):
Discrimination can generally be characterized by the morally objectionable practice of subjecting a person (or group of persons) to a treatment in some social dimension that, for no good reason, is disadvantageous compared to the treatment awarded to other persons who are in a similar situation, but who belong to another socially salient group.
A group is considered to be socially salient “if perceived membership of it is important to the structure of social interactions across a wide range of social contexts” (Weerts et al., 2023). Defining the socially salient group can only be done by agreeing on a specific, shared social context. Arriving to this point by recognizing certain groups over others under non-discrimination law captures the social phenomenon.
@@
There is a clear comparative element to this view on discrimination. We are not interested in the difference in treatment between any two individuals from different socially salient groups, but in the difference in treatment between two similar (or similarly situated) individuals from different socially salient groups. Discrimination occurs when similar (or similarly situated) individuals that differ on membership to a socially salient group are treated differently. Under this setting, discrimination becomes the opposite of equality (Weerts et al., 2023). Also under this setting, the comparator becomes a manifestation on what similar (or similarly situated) individuals look like.
@@
Let us now address proving discrimination. We consider the European Union (EU) context because, one, most fair ML works consider only the US context and, second, the EU—with the AI Act and GDPR—currently leads on the regulation front for algorithms (Álvarez et al., 2024). In doing so, we want to show the complex pipeline behind establishing discrimination. Although the discussion below concerns EU non-discrimination law, the modeling problem formalized in Section 2.1 is shared among most Western conceptions of justice (Westen, 1982) and it is, thus, a robust problem to the legal context.
There are two forms of discrimination conceived by EU non-discrimination law: direct and indirect discrimination. Under direct discrimination, the decision-maker uses the protected attribute for the decision, while under indirect discrimination the decision-maker uses a supposedly neutral attribute that is informative of the protected attribute. [^9] In US non-discrimination law, direct discrimination is refer to as disparate treatment and indirect discrimination as disparate impact (Barocas and Selbst, 2016). The EU context almost fully translates onto the US context, especially when it comes to direct discrimination and disparate treatment. For our purposes, the goal of substantive equality over formal equality is specific to the EU (Wachter et al., 2020). We come back to this point in Section 3, but, regarding the overall problem of the comparator, we can generalize across the two legal contexts. The same seems to hold for UK non-discrimination law (Sargeant and Magnusson, 2024). There are four elements in a discrimination claim, which we list below in order (Weerts et al., 2023). We add in italics the implications to modern decision flows.
@@
  • …“a protected characteristic.” The claim must be based on an attribute protected by non-discrimination law, such as gender or race. These are attributes that summarize historical processes as they were once used, or are still in use, to not only classify individuals but to purposely shape their social and economic opportunities (Sen and Wasow, 2016; Bonilla-Silva, 1997; Michaels and Reed, 2020). Non-discrimination law acknowledges the historical injustices toward members of these groups by encoding them as “protected”. That is why a discrimination claim is specific to non-discrimination law, meaning it cannot be apply to any characteristic of an individual, only to a protected one. This element, on top of the current discussions on intersectional versus multiple discrimination (Xenidis, 2020), is ongoing as there is a fear that models are creating new protected attributes.
@@
The third element is the one that concerns the comparator as it requires the use of the discrimination testing tools for obtaining the prima facie evidence. Proving discrimination, similar to describing discrimination, has a clear comparative element; here, though, it also has a modeling angle because the goal of obtaining evidence is explicit. The idea is still to compare similar (or similarly situated) individuals, though explicitly by defining $d$ in Definition 2.1, materializing what similarity between individuals is within our context.
Defining similarity is precisely the main challenge inherent to modeling the comparator $i^{\prime}$ with respect to the complainant $i$. The social context that describes discrimination can be and often is taken for granted when modeling the comparator as it is well understood. In Example 1.1, for example, we did not need to motivate females and males as socially salient groups. The same occurs with most fair ML papers that simply list protected attributes like race and gender to which the proposed methods apply (Ruggieri et al., 2023). However, a well-understood context still does not release us from making normative statements that are socially charge when we define similarity between individuals through $d$. We cannot escape defining $d$ either as only through $d$ we derive the comparator $i^{\prime}$ for the complainant $i$. In principle, as long as some suitable $d$ exists, the modeling problem of proving discrimination appears simple and even intuitive.
The simplicity, however, breaks down precisely because we must first define similarity between individuals to then obtain similar individuals and, in turn, prove discrimination. In The Empty Idea of Equality, Westen (1982) points out at this limitation of non-discrimination law by arguing that equality is a circular concept: to argue for equality we first must define what it means to be equal. It follows, thus, that similarity is a circular concept. The circularity of similarity matters because, although a discrimination tool’s main goal is to find evidence, which on paper might seem an objective goal, any discrimination testing tool—be it implicitly or explicitly—needs to implement a normative statement on similarity in the form of some $d$ in order to be operational. It is a difficulty inherent to deriving the comparator, be it the standard one or not. Hence, none of these tools can be sold as unbiased in practice.